Use the data in the Table 3.1 below to answer the following questions: Table 3.1 Death...

  1. Use the data in the Table 3.1 below to answer the following questions:

Table 3.1 Death rates from lung cancer, per 1,000 people aged 35 or older/year

Death rates per 1000

Cigarette smokers

0.96

Nonsmokers

0.07

  1. Compute the relative risk of smokers versus nonsmokers

  1. Explain your answer to (A) above in narrative form.

  1. Compute the risk difference for smokers (Note: the rates are already expressed per 1000 so you do not need to multiply or divide your answer per 1000. Your answer can simply be stated per 1000).

  1. Explain your answer to (C) above in narrative form.
Answer #1

SOLUTION:

A.
The relative risk of smokers versus nonsmokers = 0.96/0.07 = 13.71


B.

The risk of dying in 35 + year old peoples increases 13.71 times in ease of cigarette smokers as compared to non-smokers.


C.
The risk difference for smokers = 0.96 – 0.07 = 0.89